OpenStudy (ksaimouli):

3cosx+9=7 slove for x

myininaya (myininaya):

Subtract 9 on both sides Then divide both sides by 3

OpenStudy (ksaimouli):

yes i got 131.8 another on ew

OpenStudy (ksaimouli):

plz say plz

myininaya (myininaya):

\[\cos(x)=\frac{-2}{3}\] \[\cos(x)=\cos(x+2 \pi n)\] \[\text{ so we have } \cos(x+2 n \pi)=\frac{-2}{3}\]

myininaya (myininaya):

Now take arccos of both sides

myininaya (myininaya):

\[x+2 n \pi=arc \cos(-\frac{2}{3})\]

myininaya (myininaya):

to solve for x subtract 2npi on both sides

OpenStudy (karatechopper):

@myin on this..i have to learn this for ACT, but how would u solve it exactly? what does cos mean?

myininaya (myininaya):